OpenStudy (mathisfun13):

Find the value of S for an infinite geometric series with r=1/6 and a1=10

OpenStudy (anonymous):

use \[\frac{a}{1-r}\] with \[a=10, r=\frac{1}{2}\]

OpenStudy (anonymous):

ooops i meant \(r=\frac{1}{6}\)

OpenStudy (anonymous):

you can almost do it in your head one minus one sixth is five sixths and ten divided by five sixths is ten times six fifths, i.e. \[10\times \frac{6}{5}\] then five goes in to ten twice, and two times six is twelve